please help me find the y intercept and x intercept for this equation; 16y-4x=-8
Found 3 solutions by scott8148, rfer, RAY100:
Answer by scott8148(6628)

when the graph crosses the y-axis (y-intercept), the value of x is zero
___ so 16y = -8 and y = -1/2, making the intercept (0, -1/2)

similarly, when the graph crosses the x-axis (x-intercept), the value of y is zero
___ so -4x = -8 and x = 2, making the intercept (2, 0)

Answer by rfer(16322)

y int is when x=0 (0,-1/2)
x int is when y=0 (2,0)
16y-4x=-8
16y-4(0)=-8
y=-1/2
----------------
16(0)-4x=-8
x=2

Answer by RAY100(1637)

16y -4x =-8
.
x intercept is @y=0,,,,,16(0) -4x =-8,,,,x=2
.
y intercept is @ x=0,,,,,16y -4(0) =-8,,,,,y= - 8/16 = -1/2
.
.
check,,,,16y -4x =-8,,,,,16y = 4x -8,,,,y= 1/4 *x - 1/2,,,,yintercept = - 1/2
.
.
(2,0) ,,,,,16*0 -4(2) =-8,,,ok
.
(0,-1/2),,,,,,16*-1/2 -4(0) = -8,,,,,ok
.